# Causality of higher-spin interactions on the (A)dS lightcone, with application to the static patch

### v1

@@ I.1 Scope and goals
Higher-Spin (HS) Gravity Vasiliev:1990en ; Vasiliev:1995dn ; Vasiliev:1999ba is the putative interacting theory of an infinite tower of massless gauge fields with increasing spin. It can be thought of as a smaller cousin of string theory, with stringy features such as an AdS/CFT holographic formulation Klebanov:2002ja ; Sezgin:2002rt ; Sezgin:2003pt ; Giombi:2012ms , but “native” to 4 spacetime dimensions. Here, we will focus on the so-called minimal type-A theory, which has a single parity-even field of every even spin $s=0,2,4,\dots$. What is remarkable about this version of the theory is that its AdS/CFT duality can be extended to de Sitter space, providing a working model of dS4/CFT3 Anninos:2011ui . The grain of salt is that this simplest version of HS theory does not have a limit where the higher-spin fields decouple, or where the graviton’s interactions are those of General Relativity (GR).
@@
    We work in a Poincare coordinate patch, in a lightcone gauge. The initial data on the past horizon is then expressed as data on “past lightlike infinity” in these Poincare coordinates, or, equivalently, as massless plane-wave modes.
@@
In the present paper, we aim to extend this method to “true” higher-spin interactions, i.e. interactions with arbitrary numbers of derivatives, acting on fields with arbitrary spin. We will focus on the lowest order in interactions, i.e. on the cubic vertices, which are well-known in several formalisms – see Sleight:2016dba for the vertices in the language of Fronsdal fields Fronsdal:1978rb ; Fronsdal:1978vb , or Gelfond:2018vmi ; Didenko:2018fgx ; Didenko:2019xzz ; Gelfond:2019tac for their derivation from Vasiliev’s equations Vasiliev:1990en ; Vasiliev:1995dn ; Vasiliev:1999ba . In this paper, we will instead use the lightcone formulation, given in Ponomarev:2016lrm ; Skvortsov:2018uru , using the framework developed in Bengtsson:1983pd ; Fradkin:1991iy ; Metsaev:1999ui ; Metsaev:2003cu ; Metsaev:2018xip . This is in keeping with the strategy outlined above for the static-patch scattering problem, but with an important difference: in our work on lower-spin interactions Albrychiewicz:2021ndv ; Neiman:2023bkq ; Neiman:2024vit ; Lang:2025rxt , we started from a covariant formulation, and then adopted a lightcone gauge. In this paper, we’ll skip the covariant formulation, and start directly from a lightcone formalism. On this path, two difficulties stand out:
@@
Throughout, we’ll make extensive use of the helicity/chirality structure of massless fields in 4d and their interaction vertices. In particular, the cubic vertices in the lightcone formalism decompose into two sectors – chiral and anti-chiral – related by complex conjugation. By itself, each of these sectors generates a self-contained, though non-unitary, HS theory Skvortsov:2018jea ; Skvortsov:2020wtf ; Skvortsov:2022syz ; Sharapov:2022faa ; Sharapov:2022awp ; Didenko:2022qga (at $\Lambda=0$ in the lightcone formulation, the chiral theory is complete without higher-order vertices; at $\Lambda\neq 0$, this is an open question). To construct our generalized lightcone formalism, we will first analytically continue the chiral theory, then return to the real theory by adding in the anti-chiral complex conjugate.
@@ I.2 Summary and structure of the paper
The paper is structured as follows. In section II, we introduce Metsaev’s AdS lightcone formalism for cubic-level HS theory. We do this in a more covariant notation than the original paper Metsaev:2018xip , while casting the AdS isometry group as sitting inside the larger (but broken) conformal group. Some special cases where the conformal group is unbroken, i.e. free fields and Yang-Mills-like interactions, are analyzed in Appendix A. In section III.1, we focus on the chiral theory, and shift into a chiral field frame adapted to it. In section III.3, we capitalize on all this rewriting by extending the lightcone formalism to foliations with bulk lightcones, and to de Sitter space. In section III.4, we re-introduce the anti-chiral sector. In section IV, we expand on the geometric meaning of our generalized lightcone formalism. In section V, we describe the new causality properties that the generalized lightcone formalism makes apparent. As a side benefit, we point out in section V.3 how the lightcone fields can be packaged into more covariant quantities (which, at the linearized level, coincide with Weyl curvature components). In section VI, we apply all of the above to set up scattering computations in the de Sitter static patch. As an explicit example, in section VI.4 we consider gravity-like interactions with total helicity $h_{1}+h_{2}+h_{3}=2$. This includes Self-Dual GR and its HS generalization Krasnov:2021nsq , previously discussed in this context in Neiman:2023bkq ; Neiman:2024vit ; Lang:2025rxt . Section VII is devoted to discussion and outlook. In Appendix B, we patch a small gap in the literature, by mapping the cubic coupling constants between the lightcone formalism and the language of Fronsdal fields (where the couplings are fixed by holography).
@@ II.3 Cubic vertices in Minkowski and AdS
Cubic vertices of massless fields in 4d are characterized by the helicities $h_{i}$ ($i=1,2,3$) of the fields involved. HS gravity contains all the vertices allowed by gauge/spacetime symmetry, with the exception of the purely-scalar vertex $h_{1}=h_{2}=h_{3}=0$. All other vertices can be divided into two classes: chiral $h_{1}+h_{2}+h_{3}>0$ and anti-chiral $h_{1}+h_{2}+h_{3}<0$. Since these are related by complex conjugation, we will write out only the chiral sector explictily. To specify the conformal generators in the form (11)-(14), we will need the cubic pieces of the Lagrangian $\mathcal{L}$, the intrinsic Lorentz rotations $\mathcal{M}^{\mu\nu}$, and the special conformals $\mathcal{R}^{\mu}$. Fortunately, an immediate simplification occurs: the cubic piece of $\mathcal{M}^{\mu\nu}$ is always along the left-handed bivector $q^{\alpha}q^{\beta}$ (for chiral vertices) or the right-handed bivector $\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}$ (for anti-chiral), while the cubic piece of $\mathcal{R}^{\mu}$ simply vanishes:
@@
These can be contracted with each other, or with $\bar{q}^{\dot{\alpha}}$, which produces the lightlike derivatives $\ell\cdot\partial^{(i)}$:
@@ II.3.1 Minkowski vertices
With these ingredients, the Poincare algebra (11)-(12) closes to cubic order. For $h_{1}+h_{2}+h_{3}>1$, the vertices (26)-(27) violate dilatations, since they require a dimensionful coupling constant $C_{h_{1}h_{2}h_{3}}$. In the special case $h_{1}+h_{2}+h_{3}=1$, which includes self-dual Yang-Mills and its HS generalization Krasnov:2021nsq , the coupling is dimensionless. In fact, as we show in Appendix A, in this case the entire conformal group (11)-(14) closes to cubic order.
When any “true higher-spin” vertex is switched on, i.e. if there’s any non-vanishing coupling with $h_{1}+h_{2}+h_{3}>2$, consistency at the quartic order Ponomarev:2016lrm requires that all chiral vertices (at least, all those with even spins) are included. Their couplings are then fixed as:
@@ II.3.2 AdS vertices
$\mathcal{M}_{h_{1}h_{2}h_{3}}$ $=\frac{2\mathbb{P}^{h_{1}+h_{2}+h_{3}-1}\big((h_{2}-h_{3})\beta_{1}+(h_{3}-h_{1})\beta_{2}+(h_{1}-h_{2})\beta_{3}\big)\partial_{z}}{3(h_{1}+h_{2}+h_{3})}+\dots\ ,$ (33)
@@
where $\partial^{(0)}_{\mu}$ is a derivatice acting on the explicit function of $z$, in analogy with $\partial^{(i)}_{\mu}$. The combinations $\beta_{i}\partial_{z}$ that enter the vertices (31) can now be expressed as:
@@ III.3 Analytic continuation
    With $\bar{q}^{\dot{\alpha}}$ eliminated from the formulation, the AdS symmetry is satisfied regardless of whether or not it’s equal to $q^{\alpha}\partial_{\alpha}{}^{\dot{\alpha}}z$, i.e. regarldless of the $\ell^{\mu}\partial_{\mu}z=0$ restriction!
@@ IV.2 Locally transforming between the new lightcone frames
Clearly, we can transform from any such frame to any other using de Sitter spacetime symmetries, i.e. the $t$-orthogonal components of (78)-(81). However, as noted in section III.3, there’s a potential problem – the inverse derivative in (82) along the complex spatial vector $q^{\alpha}q^{\beta}\partial_{\beta}{}^{\dot{\alpha}}t$. We will now show that one can transfrom from any lightcone frame to any other without activating this inverse derivative (and, in the chiral theory, without any inverse derivatives at all).
@@ V.1 Domains of dependence for the field equations
We argue that the answer is yes, due to the chiral nature of the cubic vertices. Indeed, the derivatives in the cubic vertex are not just tangential to the lightlike 3d “time slice” $\ell_{\mu}x^{\mu}=\mbox{const}$, but more specifically to the totally-null 2d plane $q_{\alpha}x^{\alpha\dot{\alpha}}=\mbox{const}$ (for chiral vertices), or $\bar{q}_{\dot{\alpha}}x^{\alpha\dot{\alpha}}=\mbox{const}$ (for anti-chiral vertices). In particular, as we approach the target lightray $x^{\mu}\sim\ell^{\mu}$ from the past, the derivatives in the cubic vertex become tangential to one of the totally-null planes containing this ray: either the left-handed one $q_{\alpha}x^{\alpha\dot{\alpha}}=0$, or the right-handed one $\bar{q}_{\dot{\alpha}}x^{\alpha\dot{\alpha}}=0$. Now comes a handy geometric fact: in confomally-flat spacetimes like our (A)dS, any two lightcones that share a lightray will also share the pair of totally-null planes that contain this ray. Therefore, as we approach the target lightray $x^{\mu}\sim\ell^{\mu}$ from the past, the derivatives inside the cubic vertex (no matter how many) become tangential to any lightcone that contains it. This leads to the stated causal property of the perturbative evolution.
@@ V.2 Lightcone frames with a shared lightray
As with the lightcone-preserving transformations from secton IV.2, when these generators are evaluated on the fixed ray $x^{\mu}\sim\ell^{\mu}$, two key simplifications occur. First, the interacting piece of the intrinsic Lorentz generators (82), which is along the planes $q^{\alpha}q^{\beta}$ and $\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}$, never contributes. Thus, the transformations of the fields on the fixed ray are purely kinematical. Second, whenever the free-field piece of the intrinsic Lorentz (82) is non-vanishing, the derivatives in the numerator and denominator cancel. As a result, the transformations of the fields on the fixed ray consist only of the following:
@@ V.3 From lightcone fields to covariant quantities
If we consider only chiral interactions, the whole argument can be repeated for lightcone frame transformations that preserve not a lightray (along $\ell^{\mu}$), but a totally-null left-handed plane (along $q^{\alpha}$, or equivalently $\ell^{\alpha}$). The list of generators from section V.2 will now include all the right-handed Lorentz components $J^{\dot{\alpha}\dot{\beta}}$, and $\hat{\phi}_{h}$ will end up depending only on $\ell^{\alpha}$, not on $\bar{\ell}^{\dot{\alpha}}$.
@@ A.2 Yang-Mills-like cubic interactions
Together, these subalgebras establish all the would-be conformal generators as symmetries of the action, and guarantee the correctness of almost all commutators, except $K^{2},J^{-1}]$, $K^{1},J^{-2}]$, and $J^{12},K^{-}]$. Of these, the first two are completely analogous, and the third can be reduced to them via $J^{12}=[P^{1},K^{2}]$ and the Jacobi identity. Thus, to establish conformal symmetry, we only need to check one commutator, $K^{2},J^{-1}]$, which should vanish. More speficically, since at the free-field level the conformal algebra is guaranteed, and $K^{2}$ is kinematical, we only need to check that the free-field $K^{2}$ commutes with the interaction term in $J^{-1}$:

### v2

@@ I.1 Scope and goals
Higher-Spin (HS) Gravity is the putative interacting theory of an infinite tower of massless gauge fields with increasing spin. It can be thought of as a smaller cousin of string theory, with stringy features such as an AdS/CFT holographic formulation Klebanov:2002ja ; Sezgin:2002rt ; Sezgin:2003pt ; Giombi:2012ms , but “native” to 4 spacetime dimensions. Here, we will focus on the so-called minimal type-A theory, which has a single parity-even field of every even spin $s=0,2,4,\dots$. What is remarkable about this version of the theory is that its AdS/CFT duality can be extended to de Sitter space, providing a working model of dS4/CFT3 Anninos:2011ui . The grain of salt is that this simplest version of HS theory does not have a limit where the higher-spin fields decouple, or where the graviton’s interactions are those of General Relativity (GR).
@@
    We work in a Poincare coordinate patch, in a lightcone gauge  [^1] For an application of lightcone gauge to the more standard problem of AdS boundary correlators, see Chowdhury:2024dcy .. The initial data on the past horizon is then expressed as data on “past lightlike infinity” in these Poincare coordinates, or, equivalently, as massless plane-wave modes.
@@
In the present paper, we aim to extend this method to “true” higher-spin interactions, i.e. interactions with arbitrary numbers of derivatives, acting on fields with arbitrary spin. We will focus on the lowest order in interactions, i.e. on the cubic vertices, which are well-known in several formalisms – see Sleight:2016dba for the vertices in the language of Fronsdal fields Fronsdal:1978rb ; Fronsdal:1978vb , or Gelfond:2018vmi ; Didenko:2018fgx ; Didenko:2019xzz ; Gelfond:2019tac for their derivation from Vasiliev’s equations Vasiliev:1990en ; Vasiliev:1995dn ; Vasiliev:1999ba . In this paper, we will instead use the lightcone formulation, given in Ponomarev:2016lrm ; Skvortsov:2018uru , using the framework developed in Bengtsson:1983pd ; Fradkin:1991iy ; Metsaev:1999ui ; Metsaev:2003cu ; Metsaev:2018xip . This is in keeping with the strategy outlined above for the static-patch scattering problem: a lightcone formalism naturally reduces the gauge-redundant covariant fields to their physical degrees of freedom, in a frame adapted to a static patch’s lightlike horizons, where initial/final data can be encoded. The main difference between the present paper and our previous work on lower-spin interactions Albrychiewicz:2021ndv ; Neiman:2023bkq ; Neiman:2024vit ; Lang:2025rxt is that in the latter we started from a covariant formulation, and then adopted a lightcone gauge. For general higher-spin interactions, that path remains possible, but seems difficult. Thus, in this paper, we’ll skip the covariant formulation, and instead start directly from a lightcone formalism. Having made that choice, two difficulties arise:
@@
Throughout, we’ll make extensive use of the helicity/chirality structure of massless fields in 4d and their interaction vertices. In particular, the cubic vertices in the lightcone formalism decompose into two sectors – chiral and anti-chiral – related by complex conjugation. By itself, each of these sectors generates a self-contained, though non-unitary, HS theory Skvortsov:2018jea ; Skvortsov:2020wtf ; Skvortsov:2022syz ; Sharapov:2022faa ; Sharapov:2022awp ; Didenko:2022qga (at $\Lambda=0$ in the lightcone formulation, the chiral theory is complete without higher-order vertices; at $\Lambda\neq 0$, this is an open question). Our key result is that, in the chiral sector, the generalized lightcone formalism can be read off from the original one Metsaev:2018xip via an analytic continuation. Thus, we will first restrict to the chiral theory, then analytically continue, and then return to the real theory by adding in the anti-chiral complex conjugate.
@@ I.2 Summary and structure of the paper
The paper is structured as follows. In section II, we introduce Metsaev’s AdS lightcone formalism for cubic-level HS theory. We do this in a more covariant notation than the original paper Metsaev:2018xip , while casting the AdS isometry group as sitting inside the larger (but broken) conformal group. This turns out to be useful, because some conformal generators, even though broken, nevertheless appear as building blocks inside the unbroken isometry generators. Some special cases where the conformal group is unbroken, i.e. free fields and Yang-Mills-like interactions, are analyzed in Appendix A. In section III.1, we focus on the chiral theory, and shift into a chiral field frame adapted to it. In section III.3, we capitalize on all this rewriting by extending the lightcone formalism to foliations with bulk lightcones, and to de Sitter space. In section III.4, we re-introduce the anti-chiral sector. In section IV, we expand on the geometric meaning of our generalized lightcone formalism. In section V, we describe the new causality properties that the generalized lightcone formalism makes apparent. As a side benefit, we point out in section V.3 how the lightcone fields can be packaged into more covariant quantities (which, at the linearized level, coincide with Weyl curvature components). In section VI, we apply all of the above to set up scattering computations in the de Sitter static patch. As an explicit example, in section VI.4 we consider gravity-like interactions with total helicity $h_{1}+h_{2}+h_{3}=2$. This includes Self-Dual GR and its HS generalization Krasnov:2021nsq , previously discussed in this context in Neiman:2023bkq ; Neiman:2024vit ; Lang:2025rxt . Section VII is devoted to discussion and outlook. In Appendix B, we patch a small gap in the literature, by mapping the cubic coupling constants between the lightcone formalism and the language of Fronsdal fields (where the couplings are fixed by holography).
@@ II.3 Cubic vertices in Minkowski and AdS
Cubic vertices of massless fields in 4d are characterized by the helicities $h_{i}$ ($i=1,2,3$) of the fields involved. HS gravity contains all the vertices allowed by gauge/spacetime symmetry, with the exception of the purely-scalar vertex $h_{1}=h_{2}=h_{3}=0$. All other vertices can be divided into two classes: chiral $h_{1}+h_{2}+h_{3}>0$ and anti-chiral $h_{1}+h_{2}+h_{3}<0$. Since these are related by complex conjugation, we will write out only the chiral sector explicitly. To specify the conformal generators in the form (11)-(14), we will need the cubic pieces of the Lagrangian $\mathcal{L}$, the intrinsic Lorentz rotations $\mathcal{M}^{\mu\nu}$, and the special conformals $\mathcal{R}^{\mu}$. Fortunately, an immediate simplification occurs: the cubic piece of $\mathcal{M}^{\mu\nu}$ is always along the left-handed bivector $q^{\alpha}q^{\beta}$ (for chiral vertices) or the right-handed bivector $\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}$ (for anti-chiral), while the cubic piece of $\mathcal{R}^{\mu}$ simply vanishes:
@@
The structure (22) is tightly connected to twistor formulations of chiral theories – see e.g. Adamo:2022lah ; Mason:2025pbz ; Tran:2025uad in the higher-spin context.
The derivatives (22) can be contracted with each other, or with $\bar{q}^{\dot{\alpha}}$, which produces the lightlike derivatives $\ell\cdot\partial^{(i)}$:
@@ II.3.1 Minkowski vertices
With these ingredients, the Poincare algebra (11)-(12) closes to cubic order. For $h_{1}+h_{2}+h_{3}>1$, the vertices (26)-(27) violate dilatations, since they require a dimensionful coupling constant $C_{h_{1}h_{2}h_{3}}$. In the special case $h_{1}+h_{2}+h_{3}=1$, which includes self-dual Yang-Mills and its HS generalization Ponomarev:2017nrr ; Krasnov:2021nsq , the coupling is dimensionless. In fact, as we show in Appendix A, in this case the entire conformal group (11)-(14) closes to cubic order  [^2] We thank the referee for pointing out that this claim has an existing trail in the literature MetsaevThesis ; Ponomarev:2017nrr ..
When any “true higher-spin” vertex is switched on, i.e. if there’s any non-vanishing coupling with $h_{1}+h_{2}+h_{3}>2$ (and with one of the $h_{i}$’s negative, i.e. in the covariant language the coupling can’t be constructed as just a product of linearized curvature tensors), then consistency at the quartic order Ponomarev:2016lrm ; Serrani:2025owx requires that all chiral vertices (at least, all those with even spins) are included. Their couplings are then fixed as:
@@ II.3.2 AdS vertices
$\mathcal{M}_{h_{1}h_{2}h_{3}}$ $=\frac{2\mathbb{P}^{h_{1}+h_{2}+h_{3}-1}\big((h_{2}-h_{3})\beta_{1}+(h_{3}-h_{1})\beta_{2}+(h_{1}-h_{2})\beta_{3}\big)\partial_{z}}{3(h_{1}+h_{2}+h_{3})}+\dots\ .$ (33)
Note that the appearance of $\partial_{z}$ in the flat limit (33) is an artifact of our choice (31) to always group $\partial_{z}$’s together with $\beta_{i}$’s, which will prove convenient below.
@@
where $\partial^{(0)}_{\mu}$ is a derivative acting on the explicit function of $z$, in analogy with $\partial^{(i)}_{\mu}$. The combinations $\beta_{i}\partial_{z}$ that enter the vertices (31) can now be expressed as:
@@ III.3 Analytic continuation
    With $\bar{q}^{\dot{\alpha}}$ eliminated from the formulation, the AdS symmetry is satisfied regardless of whether or not it’s equal to $q^{\alpha}\partial_{\alpha}{}^{\dot{\alpha}}z$, i.e. regardless of the $\ell^{\mu}\partial_{\mu}z=0$ restriction!
@@ IV.2 Locally transforming between the new lightcone frames
Clearly, we can transform from any such frame to any other using de Sitter spacetime symmetries, i.e. the $t$-orthogonal components of (78)-(81). However, as noted in section III.3, there’s a potential problem – the inverse derivative in (82) along the complex spatial vector $q^{\alpha}q^{\beta}\partial_{\beta}{}^{\dot{\alpha}}t$. We will now show that one can transform from any lightcone frame to any other without activating this inverse derivative (and, in the chiral theory, without any inverse derivatives at all).
@@ V.1 Domains of dependence for the field equations
We argue that the answer is yes, due to the chiral nature of the cubic vertices. Indeed, the derivatives in the cubic vertex are not just tangential to the lightlike 3d “time slice” $\ell_{\mu}x^{\mu}=\mbox{const}$, but more specifically to the totally-null 2d plane $q_{\alpha}x^{\alpha\dot{\alpha}}=\mbox{const}$ (for chiral vertices), or $\bar{q}_{\dot{\alpha}}x^{\alpha\dot{\alpha}}=\mbox{const}$ (for anti-chiral vertices). In particular, as we approach the target lightray $x^{\mu}\sim\ell^{\mu}$ from the past, the derivatives in the cubic vertex become tangential to one of the totally-null planes containing this ray: either the left-handed one $q_{\alpha}x^{\alpha\dot{\alpha}}=0$, or the right-handed one $\bar{q}_{\dot{\alpha}}x^{\alpha\dot{\alpha}}=0$. Now comes a handy geometric fact: in conformally-flat spacetimes like our (A)dS, any two lightcones that share a lightray will also share the pair of totally-null planes that contain this ray. Therefore, as we approach the target lightray $x^{\mu}\sim\ell^{\mu}$ from the past, the derivatives inside the cubic vertex (no matter how many) become tangential to any lightcone that contains it. This leads to the stated causal property of the perturbative evolution.
@@ V.2 Lightcone frames with a shared lightray
As with the lightcone-preserving transformations from section IV.2, when these generators are evaluated on the fixed ray $x^{\mu}\sim\ell^{\mu}$, two key simplifications occur. First, the interacting piece of the intrinsic Lorentz generators (82), which is along the planes $q^{\alpha}q^{\beta}$ and $\bar{q}^{\dot{\alpha}}\bar{q}^{\dot{\beta}}$, never contributes. Thus, the transformations of the fields on the fixed ray are purely kinematical. Second, whenever the free-field piece of the intrinsic Lorentz (82) is non-vanishing, the derivatives in the numerator and denominator cancel. As a result, the transformations of the fields on the fixed ray consist only of the following:
@@ V.3 From lightcone fields to covariant quantities
If we consider only chiral interactions, the whole argument can be repeated for lightcone frame transformations that preserve not a lightray (along $\ell^{\mu}$), but a totally-null left-handed plane (along $q^{\alpha}$, or equivalently $\ell^{\alpha}$). The list of generators from section V.2 will now include all the right-handed Lorentz components $J^{\dot{\alpha}\dot{\beta}}$, and $\hat{\phi}_{h}$ will end up depending only on $\ell^{\alpha}$, not on $\bar{\ell}^{\dot{\alpha}}$. Note that the result (89)-(90) does not involve approximations other than cutting off the (A)dS symmetry generators at cubic order. In particular, this means that if the chiral theory is consistent to all orders with just cubic generators (as in the flat case), then (89)-(90) also holds to all orders in the chiral coupling.
@@ A.2 Yang-Mills-like cubic interactions
Together, these subalgebras establish all the would-be conformal generators as symmetries of the action, and guarantee the correctness of almost all commutators, except $K^{2},J^{-1}]$, $K^{1},J^{-2}]$, and $J^{12},K^{-}]$. Of these, the first two are completely analogous, and the third can be reduced to them via $J^{12}=[P^{1},K^{2}]$ and the Jacobi identity. Thus, to establish conformal symmetry, we only need to check one commutator, $K^{2},J^{-1}]$, which should vanish. More specifically, since at the free-field level the conformal algebra is guaranteed, and $K^{2}$ is kinematical, we only need to check that the free-field $K^{2}$ commutes with the interaction term in $J^{-1}$: